# Critic-Guided Reinforcement Unlearning in Text-to-Image Diffusion

### v1

@@ 1 Introduction
Text-to-image diffusion models (DMs) are the backbone of modern image synthesis, yet real deployments increasingly require unlearning—removing specific concepts (styles, identities, unsafe content) without harming overall utility ho2020denoisingdiffusionprobabilisticmodels; rombach2022highresolutionimagesynthesislatent. Existing diffusion unlearning methods typically rely on parameter edits or global penalties and can either over-suppress harmless content or be evaded by adversarial phrasing  gandikota2023erasingconceptsdiffusionmodels; gong2024reliable; huang2024recelerreliableconcepterasing; tsai2024ringabellreliableconceptremoval; pham2023circumventingconcepterasuremethods. In parallel, reinforcement learning (RL) provides a flexible route to optimize non-differentiable or composite objectives by casting denoising as a sequential decision process  black2024trainingdiffusionmodelsreinforcement; fan2023dpokreinforcementlearningfinetuning; wallace2023diffusionmodelalignmentusing; clark2024directlyfinetuningdiffusionmodels. However, most RL-for-diffusion pipelines still depend on end-of-trajectory rewards  black2024trainingdiffusionmodelsreinforcement; clark2024directlyfinetuningdiffusionmodels, which produce high-variance updates and weak credit assignment across many denoising steps. For unlearning—where signals are sparse and boundary precision matters—these limitations are especially challenging  feng2025surveygenerativemodelunlearning.
We propose Critic-Guided Reinforcement Unlearning (CGRU), an RL framework that treats the reverse diffusion kernel as the policy and augments policy-gradient updates with a per-timestep critic. Concretely, we train a reward predictor on noisy latents to deliver timestep-aware signals  radford2021learningtransferablevisualmodels; a learned value function over (prompt, latent, timestep) yields advantage estimates that stabilize optimization and sharpen credit assignment. CGRU supports off-policy reuse via importance weighting for sample efficiency  10.5555/645531.656005 and integrates with standard text-to-image backbones without architectural changes. Conditioning rewards and values on intermediate latents mitigates two common failure modes in unlearning: (i) over-suppression from global penalties and (ii) instability from sparse terminal rewards.
@@
  • We cast unlearning as RL over the reverse diffusion process and introduce a timestep-aware value function to reduce variance and improve credit assignment  schulman2017proximalpolicyoptimizationalgorithms; mohamed2020montecarlogradientestimation; williams1992reinforce.
  • We train a CLIP-based classifier on noisy latents to provide per-step signals for advantage-weighted updates, localizing forgetting within the denoising trajectory  radford2021learningtransferablevisualmodels.
  • We enable off-policy reuse with importance weighting and offer a plug-and-play design—no architectural changes to standard text-to-image backbones  10.5555/645531.656005.
@@ 2 Related Works
Diffusion models learn to reverse a noise-adding process, gradually denoising latent variables into samples ho2020denoisingdiffusionprobabilisticmodels. Latent-space variants reduce cost while keeping quality rombach2022highresolutionimagesynthesislatent. As DMs are widely deployed, steering them toward desired behaviors has become a key research thread.
@@
Beyond supervised fine-tuning, reinforcement learning (RL) offers a direct way to optimize non-differentiable or composite objectives. A fundamental step is to pose denoising as a sequential decision process and optimize downstream rewards with policy gradients, as in Denoising Diffusion Policy Optimization (DDPO) black2024trainingdiffusionmodelsreinforcement. Subsequent work frames text-to-image fine-tuning with KL-regularized policy gradients (DPOK) fan2023dpokreinforcementlearningfinetuning and adapts Direct Preference Optimization to diffusion likelihoods (Diffusion-DPO) wallace2023diffusionmodelalignmentusing, with large-scale studies reporting multi-objective preference gains zhang2024improvinggflownetstexttoimagediffusion. Other directions include direct rewards backpropagation through the sampler (DRaFT) for differentiable rewards clark2024directlyfinetuningdiffusionmodels, score-based preference optimization cai2025dspodirectsemanticpreference, and extensions to discrete diffusion borso2025preferencebasedalignmentdiscretediffusion.
Standard policy-gradient tools motivate baselines/critics to reduce variance greensmith2004variance. However, most DM alignment methods use end-of-trajectory rewards, which are sparse and high-variance, highlighting the need for timestep-aware signals.
@@
Machine unlearning aims to remove specific data or concepts from a trained model for privacy, safety, or compliance. Classic foundations include data deletion and SISA training in discriminative settings ginart2019makingaiforgetyou; bourtoule2020machineunlearning. For diffusion models, concept erasure techniques target styles, identities, or unsafe content while preserving benign capabilities. Representative approaches include ESD gandikota2023erasingconceptsdiffusionmodels, lightweight erasers (Receler) huang2024recelerreliableconcepterasing, closed-form editing (RECE) gong2024reliable, and localized/gated erasure lee2025localizedconcepterasuretexttoimage.
Robust forgetting remains challenging: erased concepts can often be re-elicited via adversarial prompts tsai2024ringabellreliableconceptremoval; pham2023circumventingconcepterasuremethods; beerens2025vulnerabilityconcepterasurediffusion. Evolving surveys and benchmarks continue to clarify objectives, taxonomies, and metrics in generative unlearning feng2025surveygenerativemodelunlearning.
@@
Despite fast progress in alignment and unlearning, there are few principled methods that combine them for diffusion models. Some alignment pipelines steer models away from undesirable content using preference-based objectives, DPO-style losses, or KL-regularized updates fan2023dpokreinforcementlearningfinetuning; wallace2023diffusionmodelalignmentusing. Direct Unlearning Optimization (DUO) optimizes paired unsafe/safe preferences to remove targeted concepts while preserving utility park2025directunlearningoptimizationrobust. However, many pipelines still rely on final-step rewards or hand-made penalties. In practice, this can yield high-variance credit assignment or, conversely, over-regularization, resulting in overly broad suppression—i.e., the system may unlearn the core concept rather than shaping only its boundary.
Our work connects these threads by introducing an RL fine-tuning scheme with a per-timestep critic. In contrast to prior erasure methods that edit parameters or embeddings directly and to DUO’s global preference optimization park2025directunlearningoptimizationrobust—our approach treats the sampler as a sequential policy with timestep-aware feedback.
@@ 3.1 Diffusion Models
Diffusion models learn to reverse a noise corruption process ho2020denoisingdiffusionprobabilisticmodels. Given clean data $x_{0}$, the forward process adds Gaussian noise: $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ where $\{\beta_{t}\}_{t=1}^{T}$ is a noise schedule. The reverse process learns to denoise by predicting noise: $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma_{t}^{2}I)$ where $\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t))$ and $\epsilon_{\theta}$ is a neural network. The training objective minimizes: $\mathcal{L}=\mathbb{E}_{t,x_{0},\epsilon}[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|^{2}]$.
@@ 3.4 Policy Gradient Methods for Diffusion
Policy gradient methods optimize the objective $J(\theta)$ by leveraging the score function identity mohamed2020montecarlogradientestimation. This identity allows us to express the gradient of an expectation as an expectation of the gradient of the log-probability, enabling gradient estimation through sampling.
@@
The estimator alternates between collecting denoising trajectories and updating parameters via gradient descent. However, this approach suffers from high variance due to the sparse reward structure mohamed2020montecarlogradientestimation.
However, the score function estimator requires on-policy data, limiting optimization to one step per round of data collection. To enable multiple optimization steps, we can use off-policy data through importance sampling 10.5555/645531.656005:
@@ 4.4.2 Critic Training
The critic uses sinusoidal timestep embeddings and FiLM layers perez2017filmvisualreasoninggeneral for temporal awareness, enabling meaningful value estimates across denoising timesteps. Details are in Appendix C.
@@ 4.4.3 Policy Optimization
Policy optimization uses advantage-weighted gradients: (1) compute advantages $A_{t}=r(x_{0},c)-V_{\phi}(x_{t},c,t)$ at every timestep, (2) shuffle timesteps to break temporal dependencies, (3) apply importance weighting for off-policy data, and (4) update the policy with gradient clipping, following DDPO black2024trainingdiffusionmodelsreinforcement.
@@ 5.2 Comparison with DDPO
We compare CGRU against DDPO black2024trainingdiffusionmodelsreinforcement using the Aesthetic Score reward function. Although DDPO is not a specialized unlearning method, it serves as an ablation baseline to validate our theoretical claims regarding the superiority of dense, timestep-aware signals over sparse terminal rewards. Both methods use identical configurations, with the only difference being CGRU’s per-timestep critic versus DDPO’s terminal rewards.
@@ 5.3 Concept Removal
To validate CGRU’s effectiveness for machine unlearning, we conducted experiments following established evaluation protocols for concept removal in diffusion models zhang2024unlearncanvas. Our experimental pipeline encompasses 20 distinct object classes, enabling systematic assessment of our method’s ability to suppress specific concepts while maintaining overall generation quality.
@@
We used curated subsets from the referenced dataset kumari2023ablatingconceptstexttoimagediffusion, partitioning the data into three distinct components: (1) classifier training data for reward function development, (2) critic training trajectories for value function estimation, and (3) unlearning prompts for policy optimization.
@@ 6.1 Quantitative Results
We evaluate CGRU’s performance using established metrics from the UnlearnCanvas benchmark zhang2024unlearncanvas, computed using the ViT classifier provided by the creators of UnlearnCanvas. Our evaluation focuses on three key metrics that capture different aspects of unlearning effectiveness:
Unlearning Accuracy (UA) measures the proportion of samples generated from prompts containing the target concept that are not correctly classified by the concept detector. In-domain Retain Accuracy (IRA) quantifies the model’s ability to correctly generate and classify samples containing other concepts from the same domain, ensuring that unlearning does not severely harm unrelated capabilities. We also measure image quality through Fréchet Inception Distance (FID) heusel2018gans to ensure that unlearning does not compromise the visual fidelity of generated images.
@@
ESD gandikota2023erasingconceptsdiffusionmodels $92.15\%$ $55.78\%$ $65.55$
FMN zhang2024unlearncanvas $45.64\%$ $90.63\%$ $131.37$
UCE gandikota2024unifiedconcepteditingdiffusion $\underline{94.31}\%$ $39.35\%$ $182.01$
CA kumari2023ablatingconceptstexttoimagediffusion $46.67\%$ $90.11\%$ 54.21
SalUn fan2024salunempoweringmachineunlearning $86.91\%$ $\mathbf{96.35}\%$ $61.05$
SEOT li2024wantdontimagecontent $23.25\%$ $\underline{95.57}\%$ $62.38$
SPM lyu2024onedimensionaladapterruleall $71.25\%$ $90.79\%$ $\underline{59.79}$
EDiff wu2024erasingundesirableinfluencediffusion $86.67\%$ $94.03\%$ $81.42$
SHS wu2024scissorhandsscrubdatainfluence $80.73\%$ $81.15\%$ $119.34$
@@ 8 Conclusion
Future work should explore adaptive reward functions, investigate robustness against adversarial prompts, and extend the framework to other unlearning types. Our code and evaluation scripts are available at code-cgru to facilitate reproducibility and encourage further research in RL-based diffusion unlearning.
@@ A.2 Objective and the DDPO score-function gradient
In DDPO black2024trainingdiffusionmodelsreinforcement, the authors take advantage of the stepwise likelihoods $p_{\theta}(x_{t-1}|x_{t},c)$ and apply the score-function (likelihood-ratio) identity at each denoising step:
@@ A.4 Baseline-augmented (advantage) gradient
Let’s start from the score-function identity mohamed2020montecarlogradientestimation. For any integrable function $f(\tau)$ and policy $p_{\theta}(\tau)$,
@@ Appendix C Feature-wise Linear Modulation (FiLM) Layers
Feature-wise Linear Modulation (FiLM) layers perez2017filmvisualreasoninggeneral are a conditioning mechanism that allows neural networks to adapt their behavior based on external conditioning information. In our critic architecture, FiLM layers enable the network to modulate its intermediate representations based on the timestep information.
@@ D.4 Implementation Details
Our implementation leverages LoRA (Low-Rank Adaptation) hu2021loralowrankadaptationlarge for efficient fine-tuning of the diffusion model, significantly reducing memory requirements while maintaining training effectiveness. The gradient accumulation strategy enables effective batch sizes larger than what would fit in GPU memory, while the increased number of denoising steps (50 vs. the typical 20) provides more detailed trajectory information for the critic network.

### v2

@@ 1 Introduction
Text-to-image diffusion models (DMs) are the backbone of modern image synthesis, yet real deployments increasingly require unlearning—removing specific concepts (styles, identities, unsafe content) without harming overall utility Ho et al. (2020); Rombach et al. (2022). Existing diffusion unlearning methods typically rely on parameter edits or global penalties and can either over-suppress harmless content or be evaded by adversarial phrasing  Gandikota et al. (2023); Gong et al. (2024); Huang et al. (2024); Tsai et al. (2024); Pham et al. (2023). In parallel, reinforcement learning (RL) provides a flexible route to optimize non-differentiable or composite objectives by casting denoising as a sequential decision process  Black et al. (2024); Fan et al. (2023); Wallace et al. (2023); Clark et al. (2024). However, most RL-for-diffusion pipelines still depend on end-of-trajectory rewards  Black et al. (2024); Clark et al. (2024), which produce high-variance updates and weak credit assignment across many denoising steps. For unlearning—where signals are sparse and boundary precision matters—these limitations are especially challenging  Feng et al. (2025).
We propose Critic-Guided Reinforcement Unlearning (CGRU), an RL framework that treats the reverse diffusion kernel as the policy and augments policy-gradient updates with a per-timestep critic. Concretely, we train a reward predictor on noisy latents to deliver timestep-aware signals  Radford et al. (2021); a learned value function over (prompt, latent, timestep) yields advantage estimates that stabilize optimization and sharpen credit assignment. CGRU supports off-policy reuse via importance weighting for sample efficiency  Kakade and Langford (2002) and integrates with standard text-to-image backbones without architectural changes. Conditioning rewards and values on intermediate latents mitigates two common failure modes in unlearning: (i) over-suppression from global penalties and (ii) instability from sparse terminal rewards.
@@
  • We cast unlearning as RL over the reverse diffusion process and introduce a timestep-aware value function to reduce variance and improve credit assignment  Schulman et al. (2017); Mohamed et al. (2020); Williams (1992).
  • We train a CLIP-based classifier on noisy latents to provide per-step signals for advantage-weighted updates, localizing forgetting within the denoising trajectory  Radford et al. (2021).
  • We enable off-policy reuse with importance weighting and offer a plug-and-play design—no architectural changes to standard text-to-image backbones  Kakade and Langford (2002).
@@ 2 Related Works
Diffusion models learn to reverse a noise-adding process, gradually denoising latent variables into samples Ho et al. (2020). Latent-space variants reduce cost while keeping quality Rombach et al. (2022). As DMs are widely deployed, steering them toward desired behaviors has become a key research thread.
@@
Beyond supervised fine-tuning, reinforcement learning (RL) offers a direct way to optimize non-differentiable or composite objectives. A fundamental step is to pose denoising as a sequential decision process and optimize downstream rewards with policy gradients, as in Denoising Diffusion Policy Optimization (DDPO) Black et al. (2024). Subsequent work frames text-to-image fine-tuning with KL-regularized policy gradients (DPOK) Fan et al. (2023) and adapts Direct Preference Optimization to diffusion likelihoods (Diffusion-DPO) Wallace et al. (2023), with large-scale studies reporting multi-objective preference gains Zhang et al. (2024a). Other directions include direct rewards backpropagation through the sampler (DRaFT) for differentiable rewards Clark et al. (2024), score-based preference optimization Cai et al. (2025), and extensions to discrete diffusion Borso et al. (2025).
Standard policy-gradient tools motivate baselines/critics to reduce variance Greensmith et al. (2004). However, most DM alignment methods use end-of-trajectory rewards, which are sparse and high-variance, highlighting the need for timestep-aware signals.
@@
Machine unlearning aims to remove specific data or concepts from a trained model for privacy, safety, or compliance. Classic foundations include data deletion and SISA training in discriminative settings Ginart et al. (2019); Bourtoule et al. (2020). For diffusion models, concept erasure techniques target styles, identities, or unsafe content while preserving benign capabilities. Representative approaches include ESD Gandikota et al. (2023), lightweight erasers (Receler) Huang et al. (2024), closed-form editing (RECE) Gong et al. (2024), and localized/gated erasure Lee et al. (2025).
Robust forgetting remains challenging: erased concepts can often be re-elicited via adversarial prompts Tsai et al. (2024); Pham et al. (2023); Beerens et al. (2025). Evolving surveys and benchmarks continue to clarify objectives, taxonomies, and metrics in generative unlearning Feng et al. (2025).
@@
Despite fast progress in alignment and unlearning, there are few principled methods that combine them for diffusion models. Some alignment pipelines steer models away from undesirable content using preference-based objectives, DPO-style losses, or KL-regularized updates Fan et al. (2023); Wallace et al. (2023). Direct Unlearning Optimization (DUO) optimizes paired unsafe/safe preferences to remove targeted concepts while preserving utility Park et al. (2025). However, many pipelines still rely on final-step rewards or hand-made penalties. In practice, this can yield high-variance credit assignment or, conversely, over-regularization, resulting in overly broad suppression—i.e., the system may unlearn the core concept rather than shaping only its boundary.
Our work connects these threads by introducing an RL fine-tuning scheme with a per-timestep critic. In contrast to prior erasure methods that edit parameters or embeddings directly and to DUO’s global preference optimization Park et al. (2025)—our approach treats the sampler as a sequential policy with timestep-aware feedback.
@@ 3.1 Diffusion Models
Diffusion models learn to reverse a noise corruption process Ho et al. (2020). Given clean data $x_{0}$, the forward process adds Gaussian noise: $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ where $\{\beta_{t}\}_{t=1}^{T}$ is a noise schedule. The reverse process learns to denoise by predicting noise: $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma_{t}^{2}I)$ where $\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t))$ and $\epsilon_{\theta}$ is a neural network. The training objective minimizes: $\mathcal{L}=\mathbb{E}_{t,x_{0},\epsilon}[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|^{2}]$.
@@ 3.4 Policy Gradient Methods for Diffusion
Policy gradient methods optimize the objective $J(\theta)$ by leveraging the score function identity Mohamed et al. (2020). This identity allows us to express the gradient of an expectation as an expectation of the gradient of the log-probability, enabling gradient estimation through sampling.
@@
The estimator alternates between collecting denoising trajectories and updating parameters via gradient descent. However, this approach suffers from high variance due to the sparse reward structure Mohamed et al. (2020).
However, the score function estimator requires on-policy data, limiting optimization to one step per round of data collection. To enable multiple optimization steps, we can use off-policy data through importance sampling Kakade and Langford (2002):
@@ 4.4.2 Critic Training
The critic uses sinusoidal timestep embeddings and FiLM layers Perez et al. (2017) for temporal awareness, enabling meaningful value estimates across denoising timesteps. Details are in Appendix C.
@@ 4.4.3 Policy Optimization
Policy optimization uses advantage-weighted gradients: (1) compute advantages $A_{t}=r(x_{0},c)-V_{\phi}(x_{t},c,t)$ at every timestep, (2) shuffle timesteps to break temporal dependencies, (3) apply importance weighting for off-policy data, and (4) update the policy with gradient clipping, following DDPO Black et al. (2024).
@@ 5.2 Comparison with DDPO
We compare CGRU against DDPO Black et al. (2024) using the Aesthetic Score reward function. Although DDPO is not a specialized unlearning method, it serves as an ablation baseline to validate our theoretical claims regarding the superiority of dense, timestep-aware signals over sparse terminal rewards. Both methods use identical configurations, with the only difference being CGRU’s per-timestep critic versus DDPO’s terminal rewards.
@@ 5.3 Concept Removal
To validate CGRU’s effectiveness for machine unlearning, we conducted experiments following established evaluation protocols for concept removal in diffusion models Zhang et al. (2024b). Our experimental pipeline encompasses 20 distinct object classes, enabling systematic assessment of our method’s ability to suppress specific concepts while maintaining overall generation quality.
@@
We used curated subsets from the referenced dataset Kumari et al. (2023), partitioning the data into three distinct components: (1) classifier training data for reward function development, (2) critic training trajectories for value function estimation, and (3) unlearning prompts for policy optimization.
@@ 6.1 Quantitative Results
We evaluate CGRU’s performance using established metrics from the UnlearnCanvas benchmark Zhang et al. (2024b), computed using the ViT classifier provided by the creators of UnlearnCanvas. Our evaluation focuses on three key metrics that capture different aspects of unlearning effectiveness:
Unlearning Accuracy (UA) measures the proportion of samples generated from prompts containing the target concept that are not correctly classified by the concept detector. In-domain Retain Accuracy (IRA) quantifies the model’s ability to correctly generate and classify samples containing other concepts from the same domain, ensuring that unlearning does not severely harm unrelated capabilities. We also measure image quality through Fréchet Inception Distance (FID) Heusel et al. (2018) to ensure that unlearning does not compromise the visual fidelity of generated images.
@@
ESD Gandikota et al. (2023) $92.15\%$ $55.78\%$ $65.55$
FMN Zhang et al. (2024b) $45.64\%$ $90.63\%$ $131.37$
UCE Gandikota et al. (2024) $\underline{94.31}\%$ $39.35\%$ $182.01$
CA Kumari et al. (2023) $46.67\%$ $90.11\%$ 54.21
SalUn Fan et al. (2024) $86.91\%$ $\mathbf{96.35}\%$ $61.05$
SEOT Li et al. (2024) $23.25\%$ $\underline{95.57}\%$ $62.38$
SPM Lyu et al. (2024) $71.25\%$ $90.79\%$ $\underline{59.79}$
EDiff Wu et al. (2024) $86.67\%$ $94.03\%$ $81.42$
SHS Wu and Harandi (2024) $80.73\%$ $81.15\%$ $119.34$
@@ 8 Conclusion
Future work should explore adaptive reward functions, investigate robustness against adversarial prompts, and extend the framework to other unlearning types. Our code and evaluation scripts are available at Anonymous (2025) to facilitate reproducibility and encourage further research in RL-based diffusion unlearning.
@@ A.2 Objective and the DDPO score-function gradient
In DDPO Black et al. (2024), the authors take advantage of the stepwise likelihoods $p_{\theta}(x_{t-1}|x_{t},c)$ and apply the score-function (likelihood-ratio) identity at each denoising step:
@@ A.4 Baseline-augmented (advantage) gradient
Let’s start from the score-function identity Mohamed et al. (2020). For any integrable function $f(\tau)$ and policy $p_{\theta}(\tau)$,
@@ Appendix C Feature-wise Linear Modulation (FiLM) Layers
Feature-wise Linear Modulation (FiLM) layers Perez et al. (2017) are a conditioning mechanism that allows neural networks to adapt their behavior based on external conditioning information. In our critic architecture, FiLM layers enable the network to modulate its intermediate representations based on the timestep information.
@@ D.4 Implementation Details
Our implementation leverages LoRA (Low-Rank Adaptation) Hu et al. (2021) for efficient fine-tuning of the diffusion model, significantly reducing memory requirements while maintaining training effectiveness. The gradient accumulation strategy enables effective batch sizes larger than what would fit in GPU memory, while the increased number of denoising steps (50 vs. the typical 20) provides more detailed trajectory information for the critic network.